### v1

# Hao Sun, Weiwei Deng, Furu Wei, Qi Zhang Microsoft Corporation

Abstract
The remarkable capability of large language models (LLMs) for in-context learning (ICL) needs to be activated by demonstration examples. Prior work has extensively explored the selection of examples for ICL, predominantly following the "select then organize" paradigm, such approaches often neglect the internal relationships between examples and exist an inconsistency between the training and inference. In this paper, we formulate the problem as a sequential selection problem and introduce $\texttt{Se}^{2}$, a sequential-aware method that leverages the LLM’s feedback on varying context, aiding in capturing inter-relationships and sequential information among examples, significantly enriching the contextuality and relevance of ICL prompts. Meanwhile, we utilize beam search to seek and construct example sequences, enhancing both quality and diversity. Extensive experiments across 23 NLP tasks from 8 distinct categories illustrate that $\texttt{Se}^{2}$ markedly surpasses competitive baselines and achieves 42% relative improvement over random selection. Further in-depth analysis show the effectiveness of proposed strategies, highlighting $\texttt{Se}^{2}$’s exceptional stability and adaptability across various scenarios. Our code will be released to facilitate future research.
[…]
$\texttt{Se}^{2}$: Sequential Example Selection for In-Context Learning
Haoyu Liu, Jianfeng Liu, Shaohan Huang, Yuefeng Zhan, Hao Sun, Weiwei Deng, Furu Wei, Qi Zhang Microsoft Corporation implhy@gmail.com {jianfengliu, shaohanh, yuefzh, hasun, dedeng, fuwei, qizhang}@microsoft.com
1 Introduction
Large Language Models (LLMs) Radford et al. (2019); Brown et al. (2020); Zhang et al. ; Touvron et al. (2023a) have demonstrated remarkable capabilities in handling a wide range of problems and tasks through in-context learning (ICL). ICL allows these models to learn from a limited number of examples without the need for parameter updates. Despite Wei et al. (2022c); Wang et al. (2022); Wei et al. (2022b) shows the ICL superior potential for various scenarios, the performance of ICL heavily relies on the carefully example selection Liu et al. (2021); Rubin et al. (2021); Su et al. (2022); Zhang et al. (2022); Cheng et al. (2023). Recent work Lu et al. (2022); Zhang et al. (2022) have also shown there are significant variance in performance across different sets of examples. Thus, the example selection plays a vital role in the behavior of ICL.
Figure 1: The "select then organize" paradigm and ours sequential example selection.
Given the diverse nature of NLP tasks, researchers have proposed various heuristic criteria for example selection, including entropy Lu et al. (2022); Wu et al. (2023), influences Nguyen and Wong (2023), uncertainty Diao et al. (2023), and so on. Early approaches focused on retrieving examples based on similarity between the input and the examples Liu et al. (2022); Su et al. (2022). More recent efforts Rubin et al. (2021); Cheng et al. (2023) have explored selecting examples based on their contribution to output performance. The above methods select examples and concatenate them together, following the "select then organize" paradigm, as shown in Figure 1. However, these methods often overlook the impact of the relationship between examples. To address this, Zhang et al. (2022) proposing a reinforcement learning (RL) approach to formalize the process. Yet, this approach faces limitations in modeling and applicability due to the constraints of the RL framework.
Selecting an optimal examples sequence from a large candidate pool is an NP-hard problem. To overcome this challenge, we propose a novel method for example selection, called  $\texttt{Se}^{2}$, which selects the ideal sequence by modeling the conditional probability of the example sequence given the current context.  $\texttt{Se}^{2}$ employs a sequential-aware model that utilizes feedback from a LLM across a diverse set of prompting examples as training data. This allows  $\texttt{Se}^{2}$ to model the interrelationships and sequential information between examples. During inference, we adopt a beam search strategy to construct example sequences, which can enhance the quality and diversity of prompts.
We empirically validated the effectiveness of  $\texttt{Se}^{2}$ at 23 popular tasks, including question answering, paraphrase detection and so on. $\texttt{Se}^{2}$ outperformed competitive baselines and achieved a $42\%$ relative improvement over random selection. Our quantitative evaluations demonstrate the advantages of $\texttt{Se}^{2}$, showing a significant performance boost with little variance from the sequential training pattern and improved example sequence quality through beam search. Moreover, $\texttt{Se}^{2}$ exhibits strong transferability, enhancing larger LLMs with feedback from smaller scoring LLMs. Case study also reveals  $\texttt{Se}^{2}$’s ability to identify example sequences with inherent logical relationships.
Overall, our contributions are summarized as follows:
  • In this paper, we explore a novel sequential example selection paradigm for ICL and figure out the importance of a sequential approach in the selection process.
  • We propose  $\texttt{Se}^{2}$, a sequence-aware method that can adeptly handle sequential relationships, generating ideal example sequences as in-context prompt.
  • Through extensive experimentation on 23 popular benchmarks,  $\texttt{Se}^{2}$ demonstrates significant performance improvements, highlighting its ability to uncover and leverage the intrinsic connections between selected examples.
2 Preliminary
In-context learning (ICL) is a pivotal capability of Large Language Models (LLMs), enabling these models to undertake a variety of tasks by observing task prompts without updating parameters. In-context prompts typically comprise sequences of multiple examples. Formally, a $K$-shot prompt for ICL consist of $K$ examples. Given a test sample $(x_{test},y_{test})$, LLMs predicts $\hat{y}$ based on the in-context prompt and input $x_{test}$:
$\hat{y}$ $=\mathrm{LLM}(e_{K}\oplus,...,\oplus e_{1}\oplus x_{test})$ (1)
where $e_{i}=(x_{i},y_{i})_{i=1}^{K}$ represents an example consisting of an input-output pair, $K$ denotes the shot number and $\oplus$ is the concatenation operation. Our objective is to optimize the in-context prompt by seeking the ideal example sequence $\{e_{K},\dots,e_{1}\}$ in $\mathcal{E}$ for $x_{test}$ , aiming to make the LLM’s prediction $\hat{y}$ match the ground truth $y_{test}$.
Figure 2: The process of Context Sequence Construction, mainly includes sampling examples, scoring and ranking examples for varying context input, and selecting and extending example sequences.
3 Method
In this work, we depart from the traditional "select then organize" paradigm and propose a novel sequential example selection method, namely $\texttt{Se}^{2}$, to construct in-context prompts. This section outlines our method, including example scoring, context sequence construction, training, and the inference pipeline.
3.1 Example Scoring
As described in Section 2, LLM generates $\hat{y}$ conditioned on in-context prompts composed of examples. Prior work Rubin et al. (2022); Cheng et al. (2023) indicates that selecting examples solely based on semantic similarity dose not yeild optimal performance. An intuitive and more generalized approach involves scoring by the LLM itself. Our method is adaptable to various NLP tasks, encompassing both Natural Language Understanding (NLU) and Natural Language Generation (NLG). We apply the following scoring functions for them. Given a data instance $(x,y)$ and an example $e$, we measures the benefit of $e$ for $(x,y)$ by:
NLU
$\mathrm{S}_{\mathrm{NLU}}(x,y,e)=\frac{\mathrm{LH}(y\mid e\oplus x)}{\sum_{y^{{}^{\prime}}\in\mathcal{Y}}\mathrm{LH}(y^{\prime}\mid e\oplus x)},$ (2)
where the label space is $\mathcal{Y}$, $\mathrm{LH}(\cdot\mid\cdot)$ is the per-token conditional likelihood of the LLM.
NLG
$\hat{y}=\mathrm{LLM}(e\oplus x)$ (3)
$\mathrm{S}_{\mathrm{NLG}}(x,y,e)=\mathrm{metric}(y,\hat{y}),$ (4)
where $\mathrm{metric}(\cdot)$ is the task-specific metric (e,g., Rouge Lin (2004)) to compare the prediction $\hat{y}$ and ground truth $y$.
3.2 Context Sequence Construction
Previous methods Cheng et al. (2023); Rubin et al. (2022) concatenated different examples with input $x$ of data instance $(x,y)$ and scored them separately, but this process ignored the sequential information and inherent connections among examples. Additionally, it was not consistent with the $K$-shot ICL setting (when $K\neq 1$). To address this, we construct the example sequence and find reasonably supervised signals to align this setting. Specifically, as Figure 2 shown, we sample $L$ candidate examples from the example pool $\mathcal{E}$ as $\mathcal{C}$, where $L$ is the sample size. We score and rank $\mathcal{C}$ using frozen LLM. Then, we select a example $e_{c}$ from scored $\mathcal{C}$ based on its rank:
$p(\mathrm{rank})=\frac{f(\mathrm{rank})}{\Sigma_{\mathrm{rank^{\prime}}=1}^{L}f(\mathrm{rank^{\prime}})}$ (5)
$f(\mathrm{rank})=\mathrm{exp}(-\mathrm{rank})$ (6)
where $\mathrm{rank}\in[1,L]$, the higher-scoring example is more likely to be selected and the diversity of the data is preserved. We iteratively update the context input to $e_{c}\oplus x$, continuing the process of sampling, scoring, ranking and extending example sequences until the $K$-shot data are all constructed. Algorithm 1 presents the context sequence construction procedure. Through this process, we approximate the optimal example sequence incrementally and obtain the example selection signals for varying context inputs, thereby establishing the foundation for the training process.
Algorithm 1 Context Sequence Construction
0:  data instance $(x,y)$, example pool $\mathcal{E}$, shot number $K$, sample size $L$
0:  Training Data $\mathcal{D}$ for $(x,y)$
1:  for i=1 to $K$ do
2:     $\mathcal{C}\leftarrow$ Random Sample $L$ examples from $\mathcal{E}$
3:     Score and rank $\mathcal{C}$ for $(x,y)$ based on formula 2 and 4
4:     Append $\{x,y,\mathcal{C}\}$ and their scores to $\mathcal{D}$
5:     Select $e_{c}$ from $\mathcal{C}$ based on formula  5
6:     $x\leftarrow e_{c}\oplus x$
7:  end for
3.3 Training
After completing the context sequence construction, we obtained feedback from the LLM on candidate examples for varying context inputs. We then initialize two encoders $E_{e}(\cdot)$ for examples and $E_{x}(\cdot)$ for inputs. The aim is to maximize the bi-encoder score between the current input and its most effective example. Accordingly, the top-1 example with the highest score is selected as the current input’s positive example $e^{+}$. For the set of negative samples $\mathcal{N}$, we sample $B$ examples from $\mathcal{E}$ and choose the bottom-$B$ candidates as hard-negative examples to enhance the model’s discriminatory capability. The training loss is defined using InfoNCE van den Oord et al. (2019):
$\mathcal{L}(x,e^{+},\mathcal{N})=$
$-\log\frac{\mathrm{exp}(\mathrm{sim}(x,e^{+}))}{\sum_{e^{\prime}\in\{\mathcal{N}\bigcup e^{+}\}}\mathrm{exp}(\mathrm{sim}(x,e^{\prime}))}$ (7)
where scores are computed via inner product: $\mathrm{sim}(\cdot,\cdot)=E_{e}(\cdot)^{\top}E_{x}(\cdot)$. Observing the cross-shot example sequence data allows the model to discern the optimal example for varying context input.
Figure 3: The process of searching for example sequences during inference using beam search. We draw the case of $w=1$ to illustrate this briefly.
3.4 Inference
During inference, we encode the entire example pool $\mathcal{E}$ using the trained $E_{e}(\cdot)$ and index them. For a test input $x_{test}$, we identify beneficial examples—those with the highest inner product scores, as illustrated in Figure 3. However, greedy search may not be globally optimal for sequential selection. Therefore, we utilize beam search to increase the searching space. We compute the representation of $x_{test}$ with $E_{x}(\cdot)$ and search for the top-$w$ examples, where $w$ is beam size. These examples are then concatenated with the current inputs as new context sequence, aligning with the training process. This process is repeated, encoding inputs and seeking examples to maintain the $w$ best candidate sequences until each includes $K$ examples. The final predictions are generated by LLM conditioned on the example sequences, and evaluated using the corresponding metric for each task.
4 Experiments
We conducted extensive experiments on a wide range of NLP tasks. The results and analysis of the experiments demonstrate the effectiveness and advantages of our method.
4.1 Experiment Settings
Task and Dataset
We utilized a total of 23 tasks across 8 distinct categories, including Paraphrase Detection, Common Reasoning, Natural Language Inference, Story Generation, Data-to-Text Generation, Question Answering, Sentiment Analysis, and Text Summarization, based on references (Wei et al., 2022a; Cheng et al., 2023; Li et al., 2023). The prompt templates, and dataset details for different tasks are provided in Appendix A.
Implementation Details
We use GPT-Neo-2.7B Black et al. (2022) as the scoring and inference LLM for most experiments, in line with its widespread adoption in prior research Rubin et al. (2022); Cheng et al. (2023). Both encoders were initialized with "BERT-base-uncased" Devlin et al. (2019).
For computational efficiency, we randomly select up to 10k data points for each task to construct the training data and example pool while maintaining class balance in classification tasks. Training instances lacking positive $e^{+}$ were filtered out.
We configured the sample size $L=50$, the (hard) negative number $B=20$, the shot number $K=3$, and the beam size $w=3$ to balance the performance and efficiency. The candidate example sequence with the highest score within the beam is selected as the in-context prompt.
For comprehensive details on training hyperparameters and additional experimental specifics, refer to Appendix B.
Method Paraphrase Detection Common Reasoning
MRPC QQP PAWS Average COPA HellaSwag Average
Acc. F1 Acc. F1 Acc. Acc. Acc.
Zero-shot 46.6 46.0 48.4 42.1 51.8 47.0 67.0 54.5 60.8
Random 61.4 73.1 45.5 47.2 50.4 55.5 71.8 53.7 62.8
Best-of-10 69.4 81.3 63.2 53.9 55.8 64.7 74.0 54.4 64.2
BM25 58.8 70.2 55.3 55.7 49.8 58.0 68.0 54.8 61.4
SBERT 58.8 70.1 58.1 55.9 49.0 58.4 67.0 55.0 61.0
Instructor 58.8 70.0 56.4 57.4 49.1 58.3 71.0 55.2 63.1
AES 64.2 N/A 63.0 N/A 51.7 N/A N/A N/A N/A
UPRISE 76.0 84.2 77.9 74.3 49.2 72.3 72.0 54.3 63.2
$\texttt{Se}^{2}$ 77.9 85.6 79.2 75.6 58.4 75.3 76.0 54.6 65.3
Method Natural Language Inference Story Generation
MNLI-m MNLI-mm QNLI SNLI RTE Average Roc Story Roc Ending Average
Acc. Acc. Acc. Acc. Acc. Rouge-L Rouge-L
Zero-shot 35.3 36.6 50.9 35.3 34.3 38.5 5.8 3.2 4.5
Random 35.9 35.5 51.6 33.7 51.7 41.7 14.6 17.4 16.0
Best-of-10 41.3 40.2 53.2 35.0 54.5 44.8 19.8 17.4 18.6
BM25 36.8 36.8 51.9 38.0 53.4 43.4 10.7 17.5 14.1
SBERT 37.1 38.1 52.7 39.5 48.4 43.2 10.4 17.4 13.9
Instructor 39.4 40.1 53.5 40.3 49.5 44.6 11.6 17.4 14.5
AES 43.2 29.5 61.5 35.0 47.5 43.3 N/A N/A N/A
UPRISE 62.9 64.8 72.5 75.5 55.2 66.2 18.2 18.0 18.1
$\texttt{Se}^{2}$ 69.8 69.8 80.2 78.4 56.0 70.8 20.4 17.8 19.1
Method Data-to-Text Generation Question Answering
CommonGen E2E NLG Average ARC-C ARC-E OBQA Average
Rouge-L Rouge-L Acc. Acc. Acc.
Zero-shot 14.2 7.6 10.9 29.5 48.3 43.0 40.3
Random 28.8 43.1 36.0 30.8 56.9 45.3 44.3
Best-of-10 33.9 48.7 41.3 31.8 60.5 47.2 46.5
BM25 29.9 47.9 38.9 31.8 61.6 47.2 46.9
SBERT 29.9 42.6 36.3 32.0 62.9 47.2 47.4
Instructor 30.0 43.3 36.6 32.4 64.0 47.6 48.0
AES N/A N/A N/A N/A N/A N/A N/A
UPRISE 33.0 51.9 42.5 32.9 64.1 49.8 48.9
$\texttt{Se}^{2}$ 34.6 53.4 44.0 33.3 63.3 50.0 48.9
Method Sentiment Analysis Text Summarization
SST-2 SST-5 Sent140 Average AGNews Gigaword AESLC Average
Acc. Acc. Acc. Acc Rouge-L Rouge-L
Zero-shot 52.4 28.2 64.3 48.3 38.4 1.4 1.8 13.9
Random 52.9 22.5 66.2 47.2 35.4 18.1 3.1 19.9
Best-of-10 57.3 27.7 73.3 52.8 47.8 21.8 9.2 26.3
BM25 65.6 27.5 73.3 55.5 81.7 23.3 11.6 38.9
SBERT 72.5 22.6 73.5 56.2 85.0 19.4 7.8 37.4
Instructor 80.8 20.1 85.5 62.1 87.3 19.9 10.3 39.2
AES 82.8 20.9 69.1 57.6 78.2 N/A N/A N/A
UPRISE 78.8 52.6 84.4 71.9 91.4 25.8 13.4 43.5
$\texttt{Se}^{2}$ 89.0 52.7 83.8 75.2 91.6 25.8 14.0 43.8
Table 1: Main results on various tasks. We bold the best results. The column "Average" is the mean performance of the same category task. "N/A" means that the task or metric is not applicable to the corresponding method.
Beam Size Candidates PD. CR. NLI. SG. DTG. QA. SA. TS.
$w=1$ / 74.96 64.84 70.96 18.78 43.93 48.63 74.99 43.81
$w=2$ top-1 74.99 63.79 70.65 19.01 43.90 48.63 75.04 43.71
avg 74.91 64.30 70.61 19.03 43.87 48.34 75.05 43.85
$w=3$ top-1 75.34 65.30 70.84 19.10 44.00 48.87 75.20 43.81
avg 75.32 65.15 70.79 19.00 43.85 48.36 75.19 44.20
Table 2: The average performance of the 8 categories of tasks at different beam size.
Method Comparison
We compared our method with prior competitive ICL methods using GPT-Neo-2.7B as inference LLM, including:
  • Random Selection: We randomly sample demonstrations from the example pool, repeating 10 times. We report the average performance as "Random" and the best performance as "Best-of-10".
  • BM25 Robertson and Zaragoza (2009): A commonly used sparse retriever that extends TF-IDF to rank relevant examples for the test input.
  • SBERT Reimers and Gurevych (2019): A dense retriever by computing sentence embedding. Specifically, we follow Rubin et al. (2022) to take "paraphrase-mpnet-basev2" as encoder.
  • Instructor [^1] https://instructor-embedding.github.io/ Su et al. (2023): A new competitive model for computing text embeddings given instructions. It trained on 330 tasks and achieved SOTA on 70 diverse embedding tasks. We compared ours with the released instructor-base model  [^2] We use the new version checkpoint released on 2024/01/21: https://huggingface.co/hkunlp/instructor-base.
  • Active Example Selection (AES) [^3] https://github.com/ChicagoHAI/active-example-selection. GPT-2 Medium is used in the original paper. Zhang et al. (2022): An RL-based method to select examples. For fair comparison, we extended its tasks and re-trained it with GPT-Neo-2.7B.
  • UPRISE [^4] https://github.com/microsoft/LMOps/tree/main/uprise Cheng et al. (2023): A recently proposed prompt retriever for cross-task and different LLMs. We re-trained task-specific UPRISE that aligns settings for fair comparison.
Notably, the parameters of our encoders are identical to SBERT and UPRISE, so there is no extra burden on deployments and applications.
4.2 Main Results
Table 1 present the overall results of ICL methods at various benchmarks. We found that random sampling does lead to sizable gains compared to zero-shot. This demonstrates the necessity of providing examples for LLMs in downstream tasks. Compared to random sampling, $\texttt{Se}^{2}$ shows over 42% average relative improvement. BM25 and SBERT have a significant gain over random sampling. The text embedding model Instructor benefits from extensive training data and performs better than BM25 and SBERT. It shows that the appropriate demonstraion selection is beneficial. In comparison, $\texttt{Se}^{2}$ achieves an average relative improvement of 25%, but with fewer parameters and less training data. This demonstrates the effectiveness of our data construction and modeling approach.
Meanwhile, AES, UPRISE and $\texttt{Se}^{2}$ perform better due to using LLM’s feedback to find appropriate examples. Compared with UPRISE, $\texttt{Se}^{2}$ benefits from modeling varying context inputs, capturing relationships between examples, as well as the beam search strategy, showing significantly better performance on the majority of tasks. The result shows that the sequential selection is essential for $K$-shot ICL.
$\texttt{Se}^{2}$ and AES all consider the sequence formulation. However, the features and model of AES are relatively simple, making it difficult to capture sufficient semantic information, resulting in poor performance on some tasks, and its applicability is limited. In contrast, $\texttt{Se}^{2}$ achieves better performance by sequential model training and can be applied to all tasks.
Inference Model Method 1-shot 2-shot 4-shot 8-shot 16-shot
GPT2-XL BM25 51.40 51.38 52.26 52.53 53.14
UPRISE 64.29 65.83 66.07 66.31 66.23
$\texttt{Se}^{2}$ 67.19 67.66 68.46 68.26 68.57
GPT-Neo-2.7B BM25 51.69 52.62 54.22 55.24 55.27
UPRISE 66.93 66.98 66.92 67.10 66.84
$\texttt{Se}^{2}$ 69.11 69.64 69.60 69.72 69.63
LLaMA2-7B BM25 57.31 57.87 60.19 63.34 65.44
UPRISE 69.36 69.17 71.09 72.01 72.40
$\texttt{Se}^{2}$ 71.79 72.12 72.97 73.27 73.67
Table 3: $\texttt{Se}^{2}$’s average performance across 17 NLU tasks from 1-shot to 16-shot on LLMs with parameter scales 1.5B to 7B.
4.3 Analysis
In this section, we delve into evaluating the effectiveness of $\texttt{Se}^{2}$’s strategy, examining the impact of search strategy, the stability of selected examples, the transferability across models, and providing a case study for a tangible demonstration of $\texttt{Se}^{2}$’s advantages.
On the Effect of Search Strategy
In our method, we employ beam search to seek candidate demonstration example sequences, with beam size denoted as $w$. In this section, We investigate the effect of our search strategy by setting $w=[1,2,3]$, where $w=1$ corresponds to a greedy search strategy. Performance evaluations of searched candidate sequences are depicted in Table 2, with "top-1" indicating results from the highest scoring prompts among $w$ candidates, and "avg" representing the average performance across $w$ candidate prompts. We report the average performance of the 8 categories of tasks. Results demonstrate that as beam size $w$ increases, enlarging the search space, there’s a notable improvement in performance, signifying $\texttt{Se}^{2}$ found more optimal candidate example sequences. Moreover, "top-1" performance tends to be better than "avg", indicating $\texttt{Se}^{2}$’s ability to accurately identify and prioritize the most effective candidates, thereby affirming the strategy’s efficacy. Overall, $\texttt{Se}^{2}$’s approximate search strategy successfully identifies superior example sequences, enhancing downstream task performance.
On the stability of selected examples
As suggested by Lu et al. (2022); Zhao et al. (2021a), ICL performance can vary significantly, with different in-context examples causing fluctuations from near state-of-the-art (SOTA) performance to random guessing, and different orderings may lead to strong fluctuations in the performance.
(a) Hellaswag.
(b) CommonGen.
(c) ARC-E.
(d) AESLC.
Figure 4: The performance distribution of $\texttt{Se}^{2}$ and random sampling when tuning in-context prompts.
HellaSwag CommonGen ARC-E AESLC
Original 54.62 34.47 62.92 14.44
Permutation 54.55 34.34 62.59 13.99
Table 4: Average performance of original sequences and their full permutations obtained by  $\texttt{Se}^{2}$.
In this section, we explore the stability of $\texttt{Se}^{2}$ across 4 different categories of tasks, HellaSwag, Roc Story, ARC-E, and AESLC. We randomly sample 5 distinct example sets for each task, conducting a full permutation on them and recording $5*3!=30$ results. We also perform a full permutation on 3 candidate sequences obtained through $\texttt{Se}^{2}$ with $w=3$, recording $3*3!=18$ results. Figure 4 displays the performance distribution, highlighting that random demonstration selection results in substantial performance variability, aligning with prior findings Lu et al. (2022); Li and Qiu (2023). In contrast, $\texttt{Se}^{2}$ not only enhances performance but also ensures significantly greater stability, indicating its selected examples effectively support and describe the tasks. Further analysis of the average performance of the 3 original sequences against their permutations reveals $\texttt{Se}^{2}$’s original sequences outperform alternative permutations, showcasing the benefits of modeling sequential information and example relationships for ideal sequence generation without the need for organizing reconsideration.
On the transferability of $\texttt{Se}^{2}$
In above experiments, we show the results when the inference model and the scoring model are consistent. As the LLM is scaling up, aligning the scoring and inference models is time-consuming and resource-intensive. Thus, We explored $\texttt{Se}^{2}$’s example selection effectiveness across different LLMs, transferring it to 2 various size model GPT2-XL (1.5B) Radford et al. (2019) and LLAMA2-7B Touvron et al. (2023b), comparing against BM25 and UPRISE baselines. Table 3 illustrates that $\texttt{Se}^{2}$ consistently outperforms BM25 and UPRISE from 1-shot to 16-shot across models ranging from 1.5B to 7B parameters. Despite $\texttt{Se}^{2}$ being trained on 3-shot data, it shows a progressive performance increase with more examples, especially enhancing smaller models due to their greater reliance on high-quality examples for task execution, which is in line with the recent findings Wang et al. (2023).
Figure 5: An example on ARC-C where $\mathtt{Se}^{2}$ helps LLM infer the correct answer, but UPRISE does not.
Case Study
For an intuitive grasp of the $\mathtt{Se}^{2}$’s effectiveness, we present an interesting case on the ARC-C task. In Figure 5, We compare prompts derived from UPRISE and $\mathtt{Se}^{2}$, indicating the examples’ answers with underscores. UPRISE-selected examples, while relevant to science, were somewhat broad, leading to incorrect model predictions. Conversely, $\texttt{Se}^{2}$-selected examples related to gravity and energy conversion to kinetic energy, exhibiting a conceptually progressive relationship that guided the model to the correct answer. This demonstrates how $\texttt{Se}^{2}$’s sequential formulation fosters a rich, informative relationship between examples, enhancing LLM prediction accuracy.
5 Related Work
Selecting appropriate in-context examples for LLMs is crucial for enhancing performance in downstream tasks. Due to the variety of NLP tasks, various heuristic criteria have been proposed, including entropy Lu et al. (2022); Wu et al. (2023), diversity Ye et al. (2023b); Hongjin et al. (2022), influences Nguyen and Wong (2023), sensitivity Chen et al. (2023), and uncertainty Diao et al. (2023). Another research line posits that beneficial examples are those semantically similar to the test inputs, leading Agrawal et al. (2023) to adopt BM25 Robertson and Zaragoza (2009) to retrieve examples in machine translation, Liu et al. (2022); Lee et al. (2022) to retrieve example with a dense encoder Devlin et al. (2019). However, the assumption that model performance consistently correlates with semantic similarity is not universally reliable. Consequently, methods like EPR Rubin et al. (2022), CEIL Ye et al. (2023a), UDR Li et al. (2023), and UPRISE Cheng et al. (2023) utilize LLMs to score examples so as to select examples that the model truly prefers.
Despite these advancements, existing methods often overlook the internal relationships between examples, a factor that significantly influences model performance Zhang et al. (2022). Meanwhile, ICL prompts are often sequences composed of multiple examples, modeling such a sequence is a kind of NP-hard problem. Zhang et al. (2022) introduced a RL algorithm to develop generalized policies for example selection. However, this method simply utilizes the number of examples as the feature representation and MLP as the Q-network within the constraints of RL settings, struggles to capture the nuanced semantic features of natural language. This paper proposes a novel approach to the sequential example selection paradigm for ICL, demonstrating a more effective method for enhancing ICL performance than previously considered.
6 Conclusion
In this work, we explore a new sequential example selection paradigm for ICL and propose $\texttt{Se}^{2}$, a sequential-aware method that can end-to-end select ideal example sequences for test inputs. $\texttt{Se}^{2}$ leverage the feedback from LLMs across a broad spectrum of inputs and examples. This approach not only facilitates the modeling of sequential information and the intrinsic connections between examples but also empowers the construction of prompts through the beam search strategy. Through extensive experimentation, $\texttt{Se}^{2}$ demonstrated superior performance over established baselines, highlighting its ability to generate more effective prompts through beam search.
Our analysis revealed $\texttt{Se}^{2}$’s effectiveness and robustness in example selection, contributing to its stability and adaptability across different tasks and LLMs. This work underscores the importance of sequential example selection in improving ICL, offering valuable insights for future research in natural language processing.
Limitations
Our research primarily utilized GPT-Neo-2.7B Black et al. (2022) for experiments, constrained by computational resource limitations. This choice, while effective, also meant confronting the model’s sequence length limitations, potentially leading to the truncation of overly long inputs. Recent advancements, such as those by Ge et al. (2024), have explored innovative approaches for compressing inputs and prompts to expand the usable context window, presenting a valuable direction for future exploration. Furthermore, evolving from a BERT-based encoder to a more capable decoder-only model represents another promising avenue for enhancing our framework’s capacity and effectiveness. We see exploring this in our future work.
Additionally, the presence of various biases within LLMs, as identified in studies by Zhao et al. (2021b); Fei et al. (2023), poses a challenge. Since our method relies on feedback from LLMs, it’s conceivable that our results could be influenced by these inherent biases. Addressing this, we recognize the exploration of strategies to achieve fair and explainable outcomes from these complex models as an essential and promising area of research.
Our discussions aim to inspire further investigation within the community, encouraging advancements that address these limitations and propel the field of NLP research forward.
Appendices
Appendix A Overview of tasks and datasets
  • Paraphrase Detection: MRPC Dolan and Brockett (2005), QQP Wang et al. (2018), and Paws Wiki Zhang et al. (2019).
  • Commonsense Reasoning: COPA Roemmele et al. (2011), and HellaSwag Zellers et al. (2019).
  • Natural Language Inference: MNLI-m/mm Williams et al. (2018), QNLI Rajpurkar et al. (2018), SNLI Bowman et al. (2015), and RTE Bentivogli et al. (2009).
  • Story Generation: Roc Story and Roc Ending Mostafazadeh et al. (2016).
  • Data-to-Text Generation: CommonGen Lin et al. (2020), and E2E NLG Dušek et al. (2019).
  • Question Anwering: ARC-c/e Bhakthavatsalam et al. (2021), and OBQA Mihaylov et al. (2018).
  • Sentiment Analysis: SST-2/5 Socher et al. (2013), and Sentiment140 Go et al. (2009).
  • Text Summarization: AESLC Zhang and Tetreault (2019), AGNews Zhang et al. (2015), and Gigaword Napoles et al. (2012).
The detailed datasets statistical information are shown in Table 6 for reference. We mainly use instruction templates from FLAN Wei et al. (2022a) to convert task datasets into natural language instructions. Each task dataset corresponds to approximately seven templates. For SST5, ROC Story, and ROC Ending tasks which are not included in FLAN, we write a single template for each of them, readers can refer to Table 9 for details.
Appendix B HyperParameters and Implementation Details
Hyperparameter Assignment
Computing Infrastructure 8 V100-32GB GPUs
Number of epochs 6
Batch size per GPU 8
Maximum sequence length 512
Maximum learning rate 5e-6, 1e-5, 3e-5
Optimizer Adam
Adam epsilon 1e-8
Adam beta weights 0.9, 0.999
Learning rate scheduler warmup linear
Weight decay 0.0
Warmup steps 1000
Learning rate decay linear
Table 5: Hyperparameter settings
We list the overall hyperparameters in Table 5. Since we use two encoders initialized with "BERT-base-uncased" Devlin et al. (2019), the total number of parameters of  $\texttt{Se}^{2}$ is about 220M. The scoring stage takes about 7 hours per task, the training process takes about 9 hours per task. NLG tasks are much more time-consuming than the NLU tasks.
The subscripts in Section 3.3 range from $1$ to $L$, but in the actual $\mathrm{exp}(\cdot)$ calculation, they range from 00 to $L-1$. It is fine to use any other probability.
We extended the available tasks to AES and redid the task-specific training, but AES is currently limited in the tasks and metrics it can support. Specifically, AES requires the model to predict a single target token, the tasks that contain multiple tokens in the answer, such as multiple choice with textual content, and generation tasks are not supported for the time being. We report them as "N/A" in Table 1.
In Section 4.3, we found that some models do not follow the prompts as expected in the generation task, resulting in incomparable performance. For fair comparison, we report the average performance across all NLU tasks.
Category Task Train Test Metric
Paraphrase Detection MRPC 3,668 408 Accuracy
QQP 363,846 40,430 Accuracy
PAWS 49,401 8,000 Accuracy
Common Reasoning COPA 400 100 Accuracy
HellaSwag 39,905 10,042 Accuracy
Natural Language Inference MNLI-m 392,702 9,815 Accuracy
MNLI-mm 392,702 9,832 Accuracy
QNLI 104,743 5,463 Accuracy
SNLI 549,367 9,824 Accuracy
RTE 2,490 277 Accuracy
Story Generation Roc Story 87,526 9,799 Rouge-L
Roc Ending 87,906 9807 Rouge-L
Data-to-Text Generation CommonGen 67,389 4,018 Rouge-L
E2E NLG 33,525 1,847 Rouge-L
Question Answering ARC-C 1,117 1,165 Accuracy
ARC-E 2,241 2,365 Accuracy
OBQA 4,957 500 Accuracy
Sentiment Analysis SST-2 67,349 872 Accuracy
SST-5 8,534 2,210 Accuracy
Sent140 1,600,000 359 Accuracy
Text Summarization AGNews 120,000 7,600 Accuracy
Gigaword 2,044,465 730 Rouge-L
AESLC 13,181 1,750 Rouge-L
Table 6: The categories, statistics, split and evaluation metrics of each dataset.
Appendix C Settings and Performance of Different Work
In the experiments in this paper, we compare $\texttt{Se}^{2}$ with representative SOTA methods. Due to variations in the language models, tasks, instruction templates, training and testing datasets, as well as evaluation metrics used by different methods, and due to limitations in computational resources, it is hard to include all related work in the comparison. So we collected the settings and reported performance of previous work in Table 7 and Table 8 for reference only.
Method Mainly utilized LMs Shot Number
UDR Li et al. (2023) GPT-Neo-2.7B 8
CEIL Ye et al. (2023a) GPT-Neo-2.7B 50
AES Zhang et al. (2022) GPT-2-M 4
TopK+MDL Wu et al. (2023) GPT-2-XL 8
LENS Li and Qiu (2023) GPT-2-L 8 for SST-2 and AGNews, 10 for SST-5
GlobalE Lu et al. (2022) GPT-2/3 Series 2 for AGNews, 4 for others
Ours GPT-Neo-2.7B 3
Table 7: The setting of related method.
Method E2E NLG Commongen Roc Story Roc Ending
UDR Li et al. (2023) 32.6(BLEU-4) 27.1(BLEU-3) 17.6(BLEU-1) 24.7(BLEU-1)
Ours 53.4(Rouge-L) 34.6(Rouge-L) 20.4(Rouge-L) 17.8(Rouge-L)
Method COPA Hellaswag AGNews SST-2 SST-5 MRPC QNLI SNLI RTE
UDR Li et al. (2023) 72.8 / 91.5 92.4 50.5 / / 83.6 65.3
CEIL Ye et al. (2023a) / 43.2 / / 47.05 80.15 85.41 / /
AES Zhang et al. (2022) / / 70.8 81 / / / / /
TopK+MDL Wu et al. (2023) / / 87.94 91.51 40.27 / 61.43 58.77 /
LENS Li and Qiu (2023) / / 77.9 86.3 44.9 / / / /
GlobalE Lu et al. (2022) / / 78.1 80.2 43.2 / / / 51.3
Ours 76 54.6 91.6 89 52.7 77.9 80.2 78.4 56
Table 8: The performance of related method on the same task.
{xtabular*}
0.95p0.92 Task Category: Paraphrase Detection
Task: MRPC
Prompt Templates:
("Here are two sentences: {sentence1} {sentence2} Do they have the same meaning?", "{answer}"),
("Here are two sentences: {sentence1} {sentence2} Are the two sentences saying the same thing?", "{answer}"),
("{sentence1} {sentence2} Do the above sentences mean the same thing?", "{answer}"),
("{sentence1} {sentence2} Please tell me if the sentences above mean the same.", "{answer}"),
("{sentence1} {sentence2} Are these sentences conveying the same meaning?", "{answer}"),
("{sentence1} {sentence2} If the first sentence is true, is the second one also true?", "{answer}"),
("{sentence1} {sentence2} Are these two sentences paraphrases of each other?", "{answer}"),
("Do the following two sentences have the same meaning? {sentence1} {sentence2}", "{answer}"),
("Do these two sentences mean the same thing? {sentence1} {sentence2}", "{answer}"),
("Do these sentences have the same meaning? {sentence1} {sentence2}", "{answer}"),
Task: QQP
Prompt Templates:
("{question1} {question2} Would you say that these questions are the same?", "{answer}"),
("{question1} {question2} Do those questions have the same meaning?", "{answer}"),
("{question1} {question2} Are these two questions inquiring about the same information?", "{answer}"),
("{question1} {question2} Please tell me if those questions are the same.", "{answer}"),
("{question1} {question2} Are these two questions paraphrases of each other?", "{answer}"),
("First question: {question1} Second question: {question2} Are these two questions asking the same thing?", "{answer}"),
("Question 1: {question1} Question 2: {question2} Are questions 1 and 2 asking the same thing?", "{answer}"),
("Question 1: {question1} Question 2: {question2} Would the answer to these two questions be the same?", "{answer}"),
("Are the following two questions the same? {question1} {question2}", "{answer}"),
("Do these questions have the same meaning? {question1} {question2}", "{answer}"),
Task: PAWS
Prompt Templates:
("{sentence1} {sentence2} Do these sentences mean the same thing?", "{answer}"),
("{sentence1} {sentence2} Are these two sentences paraphrases of each other?", "{answer}"),
("1. {sentence1} 2. {sentence2} Are these two sentences paraphrases of each other?", "{answer}"),
("(1) {sentence1} (2) {sentence2} Do these two sentences mean the same thing?", "{answer}"),
("Sentence 1: {sentence1} Sentence 2: {sentence2} Do these two sentences convey the same information?", "{answer}"),
("Do these two sentences from wikipedia have the same meaning? {sentence1} {sentence2}", "{answer}"),
("Same meaning? {sentence1} {sentence2}", "{answer}"),
("Are these paraphrases? {sentence1} {sentence2}", "{answer}"),
("Do these mean the same? {sentence1} {sentence2}", "{answer}"),
("Please check if these have the same meaning. Answer ÿesïf they do, otherwise n̈o.̈ {sentence1} {sentence2}", "{answer}"),
Task Category: Paraphrase Detection
Task: COPA
Prompt Templates:
("{̈premise}Ẅhat is the {question}?", "{answer}"),
("Here is a premise: {̈premise}Ẅhat is the {question}?", "{answer}"),
("{̈premise}Ẅhat is the {question} of the preceding sentence?", "{answer}"),
("{̈premise}Ẅhat is a plausible {question}?", "{answer}"),
("Based on the following sentence, what is the {question}? {̈premise}"̈, "{answer}"),
("{̈premise}{̈question}: ", "{answer}"),
("What is the {question} of the following sentence? {̈premise}"̈, "{answer}"),
("Answer the following question about this sentence: {̈premise}Ẅhat is the {question}?", "{answer}"),
Task: HellaSwag
Prompt Templates:
("What happens next in this paragraph? {context}", "{answer}"),
("Continue writing the next sentence in this paragraph: {context}", "{answer}"),
("Continue writing the next sentence. {context}", "{answer}"),
("This is a test of commonsense. Complete the next sentence: {context}", "{answer}"),
("Write the next sentence in this paragraph: {context}", "{answer}"),
("How does the next paragraph end? {context}", "{answer}"),
("What most naturally follows? {context}", "{answer}"),
("What happens next? {context}", "{answer}"),
("What is the most logical next event? {context}", "{answer}"),
("Write the next sentence in the following story. {context}", "{answer}"),
Task Category: Natural Language Inference
Task: MNLI
Prompt Templates:
("Premise: {̈premise}Ḧypothesis: {̈hypothesis}D̈oes the premise entail the hypothesis? Yes, No, or Maybe?", "{answer}"),
("Premise: {̈premise}Ḧypothesis: {̈hypothesis}Ïs the hypothesis entailed by the premise? Yes, No, or Maybe?", "{answer}"),
("Here is a premise: {̈premise}Ḧere is a hypothesis: {̈hypothesis}Ïs it possible to conclude that if the premise is true, then so is the hypothesis? Yes, No, or Maybe?", "{answer}"),
("Sentence 1: {̈premise}S̈entence 2: {̈hypothesis}Ïs this second sentence entailed by the first sentence? Yes, No, or Maybe?", "{answer}"),
("Sentence 1: {̈premise}S̈entence 2: {̈hypothesis}Ïf the first sentence is true, then is the second sentence true? Yes, No, or Maybe?", "{answer}"),
("Based on the premise {̈premise},̈ can we conclude the hypothesis {̈hypothesis}ïs true? Yes, No, or Maybe?", "{answer}"),
("Premise: {̈premise}Ïf this premise is true, what does that tell us about whether it entails the hypothesis {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("Premise: {̈premise}B̈ased on this premise, is the hypothesis {̈hypothesis}ẗrue? Yes, No, or Maybe?", "{answer}"),
("If {̈premise},̈ can we conclude that {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("{̈premise}D̈oes it follow that {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
Task: QNLI
Prompt Templates:
("Does the sentence {̈sentence}änswer the question {̈question}?̈", "{answer}"),
("Does the sentence {̈sentence}p̈rovide a valid answer to the question {̈question}?̈", "{answer}"),
("Is {̈sentence}ä good answer to the question {̈question}?̈", "{answer}"),
("Does {̈sentence}c̈orrectly answer the question of {̈question}?̈", "{answer}"),
("Does {̈sentence}c̈ontain the correct answer to {̈question}?̈", "{answer}"),
("Q: {question} A: {sentence} Does the answer correctly answer the question?", "{answer}"),
("Question: {question} Answer: {sentence} Is the question answered in a satisfactory fashion?", "{answer}"),
("Question: {question} Is {sentence} a good answer to this question?", "{answer}"),
("Question: {question} Is {̈sentence}ẗhe correct answer?", "{answer}"),
Task: SNLI
Prompt Templates:
("If {̈premise},̈ does this mean that {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("If {̈premise},̈ can we conclude {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("If {̈premise},̈ does it logically follow that {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("Based on the sentence {̈premise},̈ is the sentence {̈hypothesis}ä true sentence? Yes, No, or Maybe?", "{answer}"),
("Premise: {premise} Hypothesis: {hypothesis} Can we conclude that the hypothesis is true if the premise is true? Yes, No, or Maybe?", "{answer}"),
("Premise: {premise} Hypothesis: {hypothesis} Given the premise, can we conclude the hypothesis? Yes, No, or Maybe?", "{answer}"),
("Here is a premise: {̈premise}Ḧere is a hypothesis: {̈hypothesis}.̈ Does the premise tell us whether the hypothesis is true? Yes, No, or Maybe?", "{answer}"),
("Is it possible to conclude that {̈premise}ïf {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("Is the premise {̈premise}ẗrue if {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
Task: RTE
Prompt Templates:
("{premise} Based on the paragraph above can we conclude that {̈hypothesis}?̈ Yes, No, or Maybe?", "{answer}"),
("{premise} Based on that paragraph can we conclude that this sentence is true? {hypothesis} Yes, No, or Maybe?", "{answer}"),
("{premise} Can we draw the following conclusion? {hypothesis} Yes, No, or Maybe?", "{answer}"),
("{premise} Does this next sentence follow, given the preceding text? {hypothesis} Yes, No, or Maybe?", "{answer}"),
("{premise} Can we infer the following? {hypothesis} Yes, No, or Maybe?", "{answer}"),
("Read the following paragraph and determine if the hypothesis is true: {premise} Hypothesis: {hypothesis} Yes, No, or Maybe?", "{answer}"),
("Read the text and determine if the sentence is true: {premise} Sentence: {hypothesis} Yes, No, or Maybe?", "{answer}"),
("Can we draw the following hypothesis from the context? Context: {premise} Hypothesis: {hypothesis} Yes, No, or Maybe?", "{answer}"),
("Determine if the sentence is true based on the text below: {hypothesis} {premise} Yes, No, or Maybe?", "{answer}"),
Task Category: Story Generation
Task: Roc Story
Prompt Templates:
("Beginning of the story: {question} Please continue the story: ", "{target}"),
Task: Roc Ending
Prompt Templates:
("Beginning of the story: {question} Please write the end of the story: ", "{target}"),
Task Category: Data-to-Text Generation
Task: CommonGen
Prompt Templates:
("Concepts: {concepts}. Write a sentence that includes all these words.", "{target}"),
("Keywords: {concepts}. What is a sentence that includes all these keywords?", "{target}"),
("Here are some concepts: {concepts}. What is a sentence about these concepts?", "{target}"),
("Produce a sentence which mentions all of these concepts: {concepts}.", "{target}"),
("Write a sentence about the following things: {concepts}.", "{target}"),
("Generate a sentence that includes all the following words: {concepts}.", "{target}"),
Task: E2E NLG
Prompt Templates:
("Attributes: {meaning_representation}. Produce a detailed sentence about this restaurant.", "{target}"),
("Data: {meaning_representation}. Can you generate a sentence about this data?", "{target}"),
("Data: {meaning_representation}. What is a sentence that describe this data?", "{target}"),
("Here are some keywords about a restaurant: {meaning_representation}. Write a sentence that describes the following attributes of a restaurant.", "{target}"),
("Here is some data about a restaurant: {meaning_representation}. Write a sentence that includes the following data about a restaurant.", "{target}"),
("Sentence: {meaning_representation}. Can you represent the content in this sentence in data form?", "{target}"),
("Write a sentence about a restaurant with all the following attributes: {meaning_representation}.", "{target}"),
("Write a sentence that is about a restaurant with all the following properties: {meaning_representation}.", "{target}"),
("Produce a detailed sentence about a restaurant using the following words: {meaning_representation}.", "{target}"),
("Generate a descriptive sentence about a restaurant using the following words: {meaning_representation}.", "{target}"),
Task Category: Question Answering
Task: ARC
Prompt Templates:
("{question}", "{answer}"),
("Question: {question} Answer:", "{answer}"),
("Question: {question} What is the correct answer to the question from the following choices?", "{answer}"),
("Q: {question} What is the correct answer to this question?", "{answer}"),
("What is the answer? {question}", "{answer}"),
("Answer the question {question}", "{answer}"),
("{question} Pick the answer from these options.", "{answer}"),
Task: OBQA
Prompt Templates:
("{fact} {question}", "{answer}"),
("Read this fact: {̈fact}N̈ow answer this question: {̈question}"̈, "{answer}"),
("Given the fact {̈fact},̈ what is the answer to the question or completion {̈question}"̈, "{answer}"),
("Knowing that {̈fact},̈ how would one answer {̈question}"̈, "{answer}"),
("Use evidence from the fact that {fact} to answer this question: {̈question}"̈, "{answer}"),
("Fact: {fact} Question: {question} What’s the answer?", "{answer}"),
("Use this fact to answer the question: {fact} {question}", "{answer}"),
Task Category: Sentiment Analysis
Task: SST-2
Prompt Templates:
("Review: {̈sentence}Ïs this movie review sentence negative or positive?", "{answer}"),
("Short movie review: {̈sentence}D̈id the critic thinking positively or negatively of the movie?", "{answer}"),
("Sentence from a movie review: {̈sentence}Ẅas the movie seen positively or negatively based on the preceding review?", "{answer}"),
("{̈sentence}Ḧow would the sentiment of this sentence be perceived?", "{answer}"),
("Is the sentiment of the following sentence positive or negative? {̈sentence}"̈, "{answer}"),
("What is the sentiment of the following movie review sentence? {̈sentence}"̈, "{answer}"),
("Would the following phrase be considered positive or negative? {̈sentence}"̈, "{answer}"),
("Does the following review have a positive or negative opinion of the movie? {̈sentence}"̈, "{answer}"),
Task: SST-5
Prompt Templates:
("Review: {̈sentence}.̈ It was", "{answer}"),
Task: Sentiment140
Prompt Templates:
("{text} What is the sentiment of this tweet?", "{answer}"),
("{text} How would the sentiment of this tweet be described?", "{answer}"),
("{text} Describe the sentiment embodied by this tweet.", "{answer}"),
("Tweet: {text} Predict the sentiment of this tweet.", "{answer}"),
("What is the sentiment of the following tweet? Tweet:{text}", "{answer}"),
("How would one describe the sentiment of this tweet? {text}", "{answer}"),
Task Category: Text Summarization
Task: AGNews
Prompt Templates:
("{̈text}Ẅhat is this text about? World, Sports, Business, or Technology?", "{answer}"),
("{̈text}Ẅhich topic is this article about? World, Sports, Business, or Technology?", "{answer}"),
("{̈text}Ẅhich is the best summary of this article? World, Sports, Business, or Technology?", "{answer}"),
("{̈text}Ẅhat is this text about? World, Sports, Business, or Technology?", "{answer}"),
("{̈text}Ẅhat best summarizes the content of the above article? World, Sports, Business, or Technology?", "{answer}"),
("Which is this about? {̈text}Ẅorld, Sports, Business, or Technology?", "{answer}"),
("Which is an appropriate title for this article? {̈text}Ẅorld, Sports, Business, or Technology?", "{answer}"),
("Select the topic that this about: {̈text}Ẅorld, Sports, Business, or Technology?", "{answer}"),
Task: Gigaword
Prompt Templates:
("Write a short summary for this text: {text}", "{summary}"),
("Briefly summarize this sentence: {text}", "{summary}"),
("Generate a short summary this sentence: {text}", "{summary}"),
("What is a shorter version of this: {text}", "{summary}"),
("{text} Write a brief summary in a sentence or less", "{summary}"),
("{text} What is a very short summary of the above text?", "{summary}"),
("{text} Summarize the aforementioned text in a single phrase.", "{summary}"),
("{text} Can you generate a short summary of the above paragraph?", "{summary}"),
Task: AESLC
Prompt Templates:
("What is the subject line for this email? {body}", "{subject}"),
("Write a subject line for this message: {body}", "{subject}"),
("{body} Write a subject line for this email.", "{subject}"),
("Here is an email: {body} What is a potential subject line for this email?", "{subject}"),
("{body} Propose a subject line for this email?", "{subject}"),
("This is the content of an email: {body} What was the subject line for this email?", "{subject}"),
("This is an email {body} What is the subject of this email?", "{subject}"),
("{body} Generate a subject line for this email.", "{subject}"),
Table 9: The prompt templates of tasks.